# Enhancing Perception Quality in Remote Sensing Image Compression via Invertible Neural Network

### v1

@@ 4.3 Ablation studies
Effectiveness of CGM. The CGM aims to incorporate guidance from synthetic images into the latent variables during the reconstruction of texture-rich images. To demonstrate the effectiveness of this approach. Table 1 presents the performance of our INN-RSIC with and without CGM on the testing sets of DOTA and UC-M. It is evident that the INN-RSIC with CGM consistently outperforms the model without CGM in terms of FID and LIPIS. This implies that encoding latent variables into Gaussian distributed variables conditioned on synthetic images can enhance the reconstruction performance during inverse mapping. This observation is consistent with the earlier study in [25].
Table 1: Quantitative comparison of our INN-RSIC with and without (w/o) QM or CGM.
Datasets Methods PSNR $\uparrow$ MS-SSIM $\uparrow$ FID $\downarrow$ LIPIS $\downarrow$
DOTA w/o QM 28.11 0.8955 236.8280 0.3521
w/o CGM 28.00 0.8949 235.6634 0.3496
INN-RSIC 27.57 0.8871 195.6175 0.3218
UC-M w/o QM 26.09 0.9061 174.5280 0.3600
w/o CGM 26.05 0.9055 173.9155 0.3556
INN-RSIC 26.00 0.9053 170.9994 0.3547
5 Conclusion
In this paper, we introduce INN-RSIC, a method aims at enhancing the perceptual quality of decoded RS images. This approach encodes the compression distortion from an image compression algorithm into Gaussian-distributed latent variables and recovers from the perspective of invertible transformation to substantially restore more texture details. Additionally, a CGM is introduced to split and encode the compression distortion information from the ground truth into latent variables conditioned on the synthetic image, thus improving the reconstruction of the enhanced image. Moreover, a QM is introduced to weaken the impact of data type conversions on reconstruction quality during the inference stage. Extensive experiments conducted on two widely used RS image datasets demonstrate that our INN-RSIC significantly outperforms existing traditional and learning-based image compression algorithms in terms of perception quality. Furthermore, our approach presents a new perspective for RS image compression algorithms to enhance texture reconstruction capabilities, particularly by minimizing the gap between synthetic and decoded images.
Appendix A Supplemental Material
In this supplementary material, Section A.1 evaluates the robustness of our INN-RSIC initially. Next, in Section A.2, we analyze the limitations of INN-RSIC and propose a research direction to further improve the texture details of decoded RS images. Subsequently, in Section A.3, we visualize the effectiveness of QM and CGM within our INN-RSIC. Finally, Section A.4 presents the model complexity.
A.1 Robustness evaluation
@@
The robustness of a model plays a critical role in practical deployment. Here, we employ two INN-RSIC models trained at two values of $\lambda$ to enhance the decoded images of ELIC across various compression ratios. The proposed INN-RSIC trained at $\lambda=8\times 10^{-4}$ is used to enhance the images compressed and decoded using ELIC at different $\lambda$ for the images in the testing set of DOTA, where $\lambda\in\{4,8,32,100,450\}\times 10^{-4}$. FID and LIPIS comparison results are depicted in Figure 6 (a). Comparing the FID values between the baseline and robust cases, it is observed that except for $\lambda=450\times 10^{-4}$, the proposed model trained at $\lambda=8\times 10^{-4}$ can still improve the perception quality of others.
@@
Figure 6: FID and LIPIS comparison of the baseline, robust, and our INN-RSIC. The baseline refers to testing on the decoded images of ELIC without additional improvement. The robust case indicates that our INN-RSIC trained at bpp 0.0735 (a) or bpp 0.1678 (b) is adopted to enhance the decoded images at a wide range of bpp.
Besides, our INN-RSIC trained with the corresponding $\lambda$ demonstrates the most significant improvement in perception quality in terms of FID and LIPIS, except for $\lambda=4\times 10^{-4}$. The poorer performance of our INN-RSIC at low bpp $0.0457$ (i.e., $\lambda=4\times 10^{-4}$) may be because the quality of the corresponding decoded image is poor at very low bpp, whereas the Gaussian distributed variables are conditioned on the decoded image. That is, the decoded image has limited guidance to recover stable results.
Furthermore, we investigate the robustness of our INN-RSIC model trained at high bitrate conditions in improving decoded images across different bitrate levels. The proposed INN-RSIC trained at $\lambda=32\times 10^{-4}$ is used to enhance the images compressed and decoded using ELIC at different $\lambda$ for the images in the testing set of DOTA, where $\lambda\in\{4,8,32,100,450\}\times 10^{-4}$. FID and LIPIS comparison results are illustrated in Figure 6 (b). Comparing the FID values between the baseline and robust cases, it is observed that except for $\lambda=450\times 10^{-4}$, the proposed model trained at $\lambda=32\times 10^{-4}$ can still improve perception quality across different bitrate levels. Moreover, our INN-RSIC trained with the corresponding $\lambda$ demonstrates the most significant improvement in perception quality in terms of FID and LIPIS, except for $\lambda=4\times 10^{-4}$.
To visualize the robust performance of the model, the robust case in Figures 7 (a) and (b) shows the enhanced images obtained by augmenting the decoded images of ELIC at high and low bitrates using the INN-RSIC model trained at low and high bitrates, respectively. The baseline denotes the image decoded by ELIC. From the results, it can be seen that although the decoded image of the robust case presents more artifacts and burrs compared to the ground truth and our INN-RSIC, it presents better visual characteristics and recovers more texture information compared to the baseline on the global vision.
In summary, the results indicate that both the proposed models trained at low and high bitrates exhibit effective enhancement of perception quality for decoded images across a wide range of bpp levels, highlighting the impressive robustness of our INN-RSIC in enhancing decoded images.
(a) The testing images "P0195" and "P0059"
(b) The testing images "P0015" and "P0089"
@@
A.2 Limitations
During training, we utilize decoded images as labels to guide the forward training of INN-RSIC in generating synthetic images, which are then fed into the inverse network along with Gaussian-distributed latent variables for image reconstruction. However, in the inference stage, as the ground truth cannot be obtained to feed into the forward network at the decoding end, i.e., the synthetic image cannot be derived, we resort to using labeled images instead of synthetic ones. In other words, the decoded images are fed into the inverse network along with Gaussian-distributed latent variables for enhancing the texture of the decoded images. While the texture information of the enhanced images is enriched with the assistance of Gaussian-distributed latent variables, the strictly invertible characteristic of INN inevitably affects the reconstruction quality of the enhanced image when using decoded images instead of synthetic ones.
This investigation primarily focuses on exploring the input sensitivity of our INN-RSIC model. Figure 8 showcases the images restored by INN-RSIC with synthetic images and decoded images as inputs during the inference stage. The results reveal that when synthetic images are used as input, the reconstructed image effectively recovers texture-rich details. However, when decoded images are employed as input, although some details are still recovered in the enhanced image, noticeable detail distortions emerge, particularly in texture-rich regions. To explore this issue further, we analyze the residual image between the synthetic image and the decoded image. It becomes evident that the differences contradict the invertible characteristic of INN, thereby inevitably limiting the performance of the final texture recovery.
@@ A.2 Limitations
Figure 8: Visualization of images generated by our INN-RSIC with different inputs during the inference stage under $\lambda=32\times 10^{-4}$. The reconstructed images are obtained using the synthetic images as input, while the enhanced images are derived using the decoded images as input. The residual images illustrate the difference between the synthetic images and the decoded images.
@@
In summary, the similarity between decoded and synthetic images significantly influences the recovery of texture-rich image details. Optimizing the algorithm design to ensure that the resulting synthetic and decoded images are as similar as possible will greatly enhance the decoding of detailed texture-rich RS images without requiring additional bitrates. This optimization holds the potential to improve the overall performance and fidelity of the INN-RSIC model, allowing for more accurate reconstruction of RS image.
A.3 QM and CGM evaluation
To visualize the effectiveness of QM and CGM, Figure 9 shows the enhanced images of our INN-RSIC with and without QM or CGM on the decoded images "P0088" and "P0097" by ELIC. The results show that although the enhanced images have higher perceptual quality than the decoded images when either QM or CGM is not available, our INN-RSIC provides the best perceptual quality when QM and CGM are used.
Figure 9: Visualization of the enhanced images generated by our INN-RSIC with and without (w/o) QM or CGM on the testing images "P0088" and "P0097" of DOTA. ELIC is adopted to derive the decoded images, which serve as input for our INN-RSIC for enhancement.
A.4 Model complexity
In this section, we evaluate the computational complexity of various image compression algorithms. Experiments are conducted on an Intel Silver 4214R CPU at 2.40GHz and a NVIDIA GeForce RTX 3090 Ti GPU using the PyTorch 2.0.1 framework with CUDA 11.7. The testing set of DOTA is utilized for the experiment. Table 2 presents the average results of the model complexity. From the results, it is evident that our INN-RSIC increases only 1.25M parameters, which accounts for just 2.30% of its baseline ELIC. This highlights its impressive characteristics in terms of memory storage. Additionally, concerning the decoding times of our INN-RSIC, it is observed to exhibit lower efficiency on the CPU. However, when utilizing a GPU, it demonstrates effective performance.
Table 2: Complexity comparison of several image compression algorithms. The "*" in "+*" denotes the increased complexity of INN-RSIC compared to its baseline, ELIC.
@@ A.4 Model complexity
Encoding Times (S) Decoding Times (S) Encoding Times (S) Decoding Times (S)
@@
ELIC 54.46 31.66 0.7329 0.8454 0.0402 0.0224
INN-RSIC (+*) +1.25 +31.91 / +0.4820 / +0.0222

### v2

@@ IV-C2 Effectiveness of CGM
CGM aims to incorporate guidance from synthetic images into the latent variables during the reconstruction of texture-rich images. To demonstrate the effectiveness of this approach. Table I presents the performance of the proposed INN-RSIC with and without CGM on the testing sets of DOTA and UC-M. It is evident that the INN-RSIC with CGM consistently outperforms the model without CGM in terms of FID and LPIPS. This implies that encoding latent variables into Gaussian distributed variables conditioned on synthetic images can enhance the reconstruction performance during inverse mapping. This observation is also consistent with the earlier research [30].
To visualize the effectiveness of QM and CGM, Fig. 7 shows the enhanced images of the proposed INN-RSIC with and without QM or CGM on the decoded images “P0088” and “P0097” by ELIC. The results indicate that although the enhanced images have higher perceptual quality than the decoded images when either QM or CGM is not available, the proposed INN-RSIC provides the best perceptual quality when QM and CGM are used.
Figure 7: Visualization of the enhanced images generated by the proposed INN-RSIC with and without (w/o) QM or CGM on the testing images “P0088” and “P0097” of DOTA. ELIC is adopted to derive the decoded images, which serve as input for the proposed INN-RSIC for enhancement.
IV-D Robustness Evaluation
@@
The robustness of a model plays a critical role in practical deployment. Here, we employ two INN-RSIC models trained at two values of $\lambda$ to enhance the decoded images of ELIC across various compression ratios. The proposed INN-RSIC trained at $\lambda=8\times 10^{-4}$ is used to enhance the images compressed and decoded using ELIC at different $\lambda$ for the images in the testing set of DOTA, where $\lambda\in\{4,8,32,100,450\}\times 10^{-4}$. FID and LPIPS comparison results are depicted in Fig. 8(a). Comparing the FID values between the baseline and robust cases, it is observed that except for $\lambda=450\times 10^{-4}$, the proposed model trained at $\lambda=8\times 10^{-4}$ can still improve the perception quality of others.
@@
Figure 8: FID and LPIPS comparison of the baseline, robust, and our INN-RSIC. The baseline refers to testing on the decoded images of ELIC without additional improvement. The robust case indicates that the proposed INN-RSIC trained at bpp 0.0735 (a) or bpp 0.1678 (b) is adopted to enhance the decoded images at a wide range of bpp.
Besides, the proposed INN-RSIC trained with the corresponding $\lambda$ demonstrates the most significant improvement in perception quality in terms of FID and LPIPS, except for $\lambda=4\times 10^{-4}$. The poorer performance of the proposed INN-RSIC at low bpp $0.0457$ (i.e., $\lambda=4\times 10^{-4}$) may be because the quality of the corresponding decoded image is poor at very low bpp, whereas the Gaussian distributed variables are conditioned on the decoded image. That is, the decoded image has limited guidance to recover stable results.
(a) Testing images “P0195” and “P0059”
(b) Testing images “P0015” and “P0089”
@@
Furthermore, we investigate the robustness of the proposed INN-RSIC model trained at high bitrate conditions in improving decoded images across different bitrate levels. The proposed INN-RSIC trained at $\lambda=32\times 10^{-4}$ is used to enhance the images compressed and decoded using ELIC at different $\lambda$ for the images in the testing set of DOTA, where $\lambda\in\{4,8,32,100,450\}\times 10^{-4}$. FID and LPIPS comparison results are illustrated in Fig. 8(b). Comparing the FID values between the baseline and robust cases, it is observed that except for $\lambda=450\times 10^{-4}$, the proposed model trained at $\lambda=32\times 10^{-4}$ can still improve perception quality across different bitrate levels. Moreover, the proposed INN-RSIC trained with the corresponding $\lambda$ demonstrates the most significant improvement in perception quality in terms of FID and LPIPS, except for $\lambda=4\times 10^{-4}$.
To visualize the robust performance of the model, the robust case in Figs. 9 (a) and (b) show the enhanced images obtained by augmenting the decoded images of ELIC at high and low bitrates using the INN-RSIC model trained at low and high bitrates, respectively. The baseline denotes the image decoded by ELIC. From the results, it can be observed that although the decoded image of the robust case presents more artifacts and burrs compared to the ground truth and the proposed INN-RSIC, it presents better visual characteristics and recovers more texture information compared to the baseline on the global vision.
In summary, the results indicate that both the proposed models trained at low and high bitrates exhibit effective enhancement of perception quality for decoded images across a wide range of bpp levels, highlighting the impressive robustness of the proposed INN-RSIC in enhancing decoded images.
@@
Figure 10: Visualization of images generated by the proposed INN-RSIC with different inputs during the inference stage under $\lambda=32\times 10^{-4}$. The reconstructed images are obtained using the synthetic images as input, while the enhanced images are derived using the decoded images as input. The residual images illustrate the difference between the synthetic images and the decoded images.
IV-E Discussion
During training, we utilize decoded images as labels to guide the forward training of INN-RSIC in generating synthetic images, which are then fed into the inverse network along with Gaussian-distributed latent variables for image reconstruction. However, in the inference stage, as the ground truth cannot be obtained to feed into the forward network at the decoding end, i.e., the synthetic image cannot be derived, so we resort to using labeled images instead of synthetic ones. In other words, the decoded images are fed into the inverse network along with Gaussian-distributed latent variables for enhancing the texture of the decoded images. While the texture information of the enhanced images is enriched with the assistance of Gaussian-distributed latent variables, the strictly invertible characteristic of INN inevitably affects the reconstruction quality of the enhanced image when using decoded images instead of synthetic ones.
This investigation primarily focuses on exploring the input sensitivity of the proposed INN-RSIC model. Fig. 10 showcases the images restored by INN-RSIC with synthetic images and decoded images as inputs during the inference stage. The results reveal that when synthetic images are used as input, the reconstructed image effectively recovers texture-rich details. However, when decoded images are employed as input, although some details are still recovered in the enhanced image, noticeable detail distortions emerge, particularly in texture-rich regions. To explore this issue further, we analyze the residual image between the synthetic image and the decoded image. It becomes evident that the differences contradict the invertible characteristic of INN, thereby inevitably limiting the performance of the final texture recovery.
@@ IV-E Discussion
In summary, the similarity between decoded and synthetic images significantly influences the recovery of texture-rich image details. Optimizing the algorithm design to ensure that the resulting synthetic and decoded images are as similar as possible will greatly enhance the decoding of detailed texture-rich RS images without requiring additional bitrates. This optimization holds the potential to improve the overall performance and fidelity of the INN-RSIC model, allowing for a more accurate reconstruction of RS images.
IV-F Model Complexity
To assess the computational complexity of the proposed method, we conduct the experiments on the DOTA testing set with an Intel Silver 4214R CPU running at 2.40GHz and one NVIDIA GeForce RTX 3090 Ti GPU. The average results are given in Table II.
TABLE II: Compare the parameters and computational cost among various image compression algorithms.
@@ IV-F Model Complexity
Encoding Times (s) Dec. Times (s) Enc. Times (s) Dec. Times (s)
HiFiC 55.42 148.24 7.1366 10.1388 0.4065 0.8096
InvComp 67.12 41.81 39.2739 42.0858 1.2498 3.6382
@@
ELIC 54.46 31.66 1.7934 1.6796 0.4671 0.2507
TCM 44.97 35.23 22.0959 21.0951 0.4637 0.4045
MLIC+ 82.36 116.48 38.5316 38.6344 0.3678 0.5007
IRN-RSIC 55.71 63.57 1.7934 2.1616 0.4671 0.2729
From the results, it is clear that MLIC+ has 82.36M parameters, whereas the proposed INN-RSIC comprises only 67.64% of that count. Notably, the INN-based enhancer we introduced adds only 1.25M parameters, representing a mere 2.30% increase over the baseline ELIC, which underscores its remarkable memory efficiency. In terms of floating point operations per second (FLOPs), HiFiC and MLIC+ bear a heavy computational load, with FLOP values of 148.24G and 116.48G, respectively. In contrast, the proposed INN-RSIC requires only 42.88% and 54.58% of the FLOPs of HiFiC and MLIC+, respectively. As for the decoding times, one can observe that the GPU decoding time of the proposed method is just 11.03% and 8.91% of that of the competitive TCM and MLIC+ algorithms, respectively.
To sum up, the proposed method requires less memory space, has lower computational complexity, and enjoys outstanding decoding speed on both CPU and GPU.
V Conclusion
In this paper, we present INN-RSIC, a novel method designed to enhance the perceptual quality of decoded RS images. This approach encodes the compression distortion of an image compression algorithm into Gaussian-distributed latent variables and leverages invertible transformations to substantially restore texture details. Additionally, we introduce the CCGM to separate and encode the compression distortion from the ground truth into latent variables conditioned on the synthetic image, thereby improving the reconstruction of the enhanced image. Moreover, we incorporate QM to mitigate the impact of data type conversions on reconstruction quality during inference. Extensive experiments on two widely used RS image datasets demonstrate that INN-RSIC achieves a superior balance between perceptual quality and high image fidelity compared to state-of-the-art image compression algorithms. Furthermore, our approach offers a new research perspective for enhancing the texture reconstruction capability of RS image compression algorithms by minimizing the gap between synthesized and decoded images.